# Leveraging Multimodal LLM Descriptions of Activity for Explainable Semi-Supervised Video Anomaly Detection

### v1

@@ 4.1 Quantitative Results
MemAE [12] 0.05 0.0 58.0
EVAL [36] 10.0 62.0 54.0
AnomalyRuler [46] N/A N/A 56.0
Scene-Graph[28] 19.0 64.0 60.0
@@
Scene-Graph +
MLLM-EVAD
25.0 70.0 63.0
@@
Auto-encoder [13] 0.3 2.0 61.0
Dictionary method [21] 1.6 10.0 48.0
Flow baseline [30] 11.0 52.0 51.0
FG Baseline [30] 21.0 53.0 61.0
EVAL [36] 24.3 64.5 N/A
Contextual GMM [49] 34.0 62.5 67.0
T-EVAL [37] 30.9 72.9 66.0
T-EVAL+
MLLM-EVAD
31.1 73.5 66.8
@@
Siamese[31] 41.2 78.6 87.2
Hybrid[20] 41.1 86.2 91.1
BG_agnostic [11] 65.1 66.9 92.3
Hybrid[20] ++
PCAB[34]
62.3 89.3 90.9
BG_agnostic[11]
++ PCAB[34]
66.0 64.9 92.9
SSMTL++v1 [4] 40.9 82.1 93.7
SSMTL++v2 [4] 47.8 85.2 91.6
AnomalyRuler [46] N/A N/A 89.7
EVAL[36] 68.2 87.6 86.0
T-EVAL[37] 67.5 89.7 88.0
T-EVAL ++
MLLM-EVAD
68.9 90.1 88.4
@@
Figure 2: Examples of anomalies in the ComplexVAD dataset that are detected by our method. In each box, the top row shows two frames of the anomalous video clip followed by the textual description generated by Gemma 3. The second row shows two frames from the closest matching exemplar to the anomaly. The textual description for the closest exemplar is also shown. By contrasting the anomalous description with the closest normal description, it is clear why the testing video clip was found to be anomalous.
@@
The results of our method on ComplexVAD, as well as those of EVAL [36], MemAE [12], AnomalyRuler [46], and Scene-Graph [28], using the three evaluation criteria described above, are reported in Table 2. We also provide the result for the combination of our method and Scene-Graph on the ComplexVAD dataset. Our MLLM-EVAD method outperforms the next best method (Scene-Graph) by 5, 4 and 1 percentage points in RBDC, TBDC, and frame-level evaluations, respectively. We achieve the best result by combining the Scene-Graph method with ours, reaching 25%, 70%, and 63% in RBDC, TBDC, and frame-level evaluations, respectively. AnomalyRuler, a recent MLLM-based method [46], achieves 54% frame-level AUC. However, we were not able to obtain results for the RBDC and TBDC criteria, as this method does not support spatial localization of anomalies. The MemAE method does very poorly for the two criteria that measure spatial localization. This implies that the regions of an image that MemAE predicts as anomalous are usually normal.
@@
For the experiments on Avenue and Street Scene, we combine our method with the best-performing existing exemplar-based method, namely Tracklet EVAL [37]. These datasets include many anomalies that need more fine-grained attributes (such as the direction and speed of travel for people and cars), so the high-level textual descriptions used by MLLM-EVAD are not sufficient on their own to detect all of the anomalies. We show that our textual descriptions do add an important new attribute, however, since, when combined with the more fine-grained attributes of Tracklet EVAL, the results with the MLLM textual descriptions improve accuracy for both datasets across all three evaluation criteria. As shown in Table 4, on the Avenue dataset, the Tracklet-EVAL + MLLM-EVAD method improves the state-of-the-art for both the RBDC and TBDC criteria. On Street Scene (results shown in Table 3), the state-of-the-art is surpassed for the TBDC criterion. Since the frame-level criterion does not measure spatial localization accuracy which is the focus of these datasets, we would argue that the slightly-below SOTA frame-level numbers are not representative of the true accuracy of these methods. We should note that we were not able to exactly reproduce the results reported for Tracklet EVAL [37] on the Avenue dataset, most likely due to different parameters, so we report the results we could reproduce for that method and use that version in combination with MLLM-EVAD for the results on the combined method.
@@ 4.2 Qualitative Results: Explainability
In addition to the quantitative results, we present qualitative results to demonstrate the explainability capabilities of our method. The combination of using an MLLM agent and an exemplar selection strategy enables intuitive explanations for detected anomalies by providing descriptions of both the anomaly and the most similar event from the nominal training video. In Figure 2, we demonstrate four detected anomalies from ComplexVAD along with their textual descriptions and the textual descriptions of the closest exemplars, as generated by our method. For example, one detected anomaly is described as ”The person is crouching down on the ground.” The closest matching exemplar from the nominal set is ”The person is walking across the grass.” The action of crouching deviates from the typical behavior observed in the training data. This difference highlights how our method leverages semantic understanding to identify uncommon activity, offering an interpretable explanation for the anomaly.
Another example involves an anomaly described as ”The person is holding a phone up and appearing to take a picture or video while being pushed in a large box by another person.” The closest matching exemplar from the nominal set is ”The person is walking along the sidewalk.”, which is, in fact, the most common activity involving a person in this dataset. The anomalous event, a person being pushed in a box, does not occur in the nominal video and represents an unusual and complex behavior. This example further demonstrates our method’s ability to detect rare interaction patterns that deviate from typical activities in the scene. In addition, these examples illustrate how our method also provides reasonable explanations for detected anomalies. By leveraging descriptive comparisons with nominal exemplars, our method enables enhanced and explainable video anomaly detection.
@@ 4.3 Ablation Study
We conduct two ablation studies by modifying the default settings of our method. In the first ablation study, instead of using sentence embeddings and cosine dissimilarity as the distance metric, we use the raw sentences and compute distances using BLEU [29] and METEOR [2]. In Table 5, we compare the results of cosine distance over Sentence-BERT embeddings with those obtained using BLEU and METEOR, based on the descriptions generated by our method using GPT-4o as the MLLM agent. The results show that Sentence-BERT embeddings yield slightly higher accuracy in the RBDC and frame-level criteria while BLEU and METEOR perform slightly better on the TBDC criterion. Sentence-BERT is the preferred similarity function considering its greater efficiency due to embeddings which can be saved and only require a dot product for comparison.
@@ 5 Limitations and Future Directions
While our framework shows promising results, we identify several limitations that pave the way for future research. First, our implementation relies on powerful but computationally expensive multimodal models like Gemma 3. Their high inference latency poses a challenge for real-time applications. Recent studies have shown that smaller, task-specific models can perform exceptionally well on specialized tasks [43, 35, 25, 26]. We will explore fine-tuning such models for single scenes, specifically for describing object activities and interactions.
Second, a primary contribution of our method is its semantic explainability, but quantitatively evaluating this feature is difficult due to a lack of suitable benchmarks. Current datasets with textual ground truth are designed for multi-scene or weakly-supervised tasks and do not fit the semi-supervised, single-scene VAD problem [55, 10, 53]. Therefore, a key future direction is the development of new datasets with ground-truth textual descriptions for scene-specific anomalies.
Finally, our framework currently leverages a robust traditional object detector, which is well-suited for scenarios with known object classes. To further generalize our approach, a promising future direction is the integration of open-vocabulary object detection methods [18, 6, 27]. By enabling the recognition of an open set of object categories, this advancement would significantly broaden our system’s applicability, allowing it to analyze and describe a more diverse range of anomalies in unconstrained, real-world environments.
@@ 6 Conclusions
We demonstrate that our method outperforms existing approaches on ComplexVAD which is a challenging benchmark specifically designed to test interaction-based anomalies. In addition our method also improves of the state-of-the-art results on standard datasets like Avenue and Street Scene when combined with the fine-grained method Tracklet EVAL [37]. Through ablation studies, we showed the importance of using sentence embeddings and detailed MLLM-generated descriptions. Notably, we found that Gemma 3 provided more informative descriptions than GPT-4o, contributing to higher detection accuracy.
@@
Supplementary Material
1 Compute Resources
@@ 2 Cropping Frames
$h$ $=\max\!\left(\tfrac{|y_{2}-y_{1}|}{2},\,h_{\text{min}}\right)$ (S3)
@@
3 Comparison of Gemma 3 and GPT-4o Activity Descriptions
Figure S1: The activity descriptions generated by Gemma 3 and GPT-4o with given images.
In Figure S1, we share activity descriptions generated by Gemma 3 and GPT-4o for given images and the promt: ”Briefly describe what the <object name>s in the enclosed regions of these images are doing. The two images were taken one second apart.”

### v2

@@ 4.1 Quantitative Results
MemAE (Gong et al., 2019) 0.05 0.0 58.0
EVAL (Singh et al., 2023) 10.0 62.0 54.0
AnomalyRuler (Wu et al., 2024b) N/A N/A 56.0
Scene-Graph (Mumcu et al., 2025b) 19.0 64.0 60.0
@@
Scene-Graph + MLLM-EVAD 25.0 70.0 63.0
@@
Auto-encoder (Hasan et al., 2016) 0.3 2.0 61.0
Dictionary method (Lu et al., 2013) 1.6 10.0 48.0
Flow baseline (Ramachandra and Jones, 2020) 11.0 52.0 51.0
FG Baseline (Ramachandra and Jones, 2020) 21.0 53.0 61.0
EVAL (Singh et al., 2023) 24.3 64.5 N/A
Contextual GMM (Yang and Radke, 2025) 34.0 62.5 67.0
T-EVAL (Singh et al., 2024) 30.9 72.9 66.0
T-EVAL + MLLM-EVAD 31.1 73.5 66.8
@@
Siamese AE (Ramachandra et al., 2020b) 41.2 78.6 87.2
Hybrid AE (Liu et al., 2021) 41.1 86.2 91.1
BG-Agnostic (Georgescu et al., 2021) 65.1 66.9 92.3
Hybrid AE (Liu et al., 2021) + PCAB (Ristea et al., 2021) 62.3 89.3 90.9
BG-Agnostic (Georgescu et al., 2021) + PCAB (Ristea et al., 2021) 66.0 64.9 92.9
SSMTL++v1 (Barbalau et al., 2023) 40.9 82.1 93.7
SSMTL++v2 (Barbalau et al., 2023) 47.8 85.2 91.6
AnomalyRuler (Wu et al., 2024b) N/A N/A 89.7
EVAL (Singh et al., 2023) 68.2 87.6 86.0
T-EVAL (Singh et al., 2024) 67.5 89.7 88.0
T-EVAL + MLLM-EVAD 68.9 90.1 88.4
@@
Figure 2: Examples of anomalies in the ComplexVAD dataset that are detected by our method. In each box, the top row shows two frames of the anomalous video clip followed by the textual description generated by Gemma 3. The second row shows two frames from the closest matching exemplar to the anomaly. The textual description for the closest exemplar is also shown. By contrasting the anomalous description with the closest normal description, it is clear why the testing video clip was found to be anomalous. The nominal description shown corresponds to the closest exemplar in the learned normal set and is presented for contrast; the anomaly is detected because its description is dissimilar from all nominal exemplars, and this semantic distance determines the anomaly score.
@@
The results of our method on ComplexVAD, as well as those of EVAL (Singh et al., 2023), MemAE (Gong et al., 2019), AnomalyRuler (Wu et al., 2024b), and Scene-Graph Mumcu et al. (2025b), using the three evaluation criteria described above, are reported in Table 2. We also provide the result for the combination of our method and Scene-Graph on the ComplexVAD dataset. Our MLLM-EVAD method outperforms the next best method (Scene-Graph) by 5, 4 and 1 percentage points in RBDC, TBDC, and frame-level evaluations, respectively. We achieve the best result by combining the Scene-Graph method with ours, reaching 25%, 70%, and 63% in RBDC, TBDC, and frame-level evaluations, respectively. AnomalyRuler, a recent MLLM-based method (Wu et al., 2024b), achieves 54% frame-level AUC. However, we were not able to obtain results for the RBDC and TBDC criteria, as this method does not support spatial localization of anomalies. The MemAE method does very poorly for the two criteria that measure spatial localization. This implies that the regions of an image that MemAE predicts as anomalous are usually normal.
@@
For the experiments on Avenue and Street Scene, we combine our method with the best-performing existing exemplar-based method, namely Tracklet EVAL (Singh et al., 2024). These datasets include many anomalies that need more fine-grained attributes (such as the direction and speed of travel for people and cars), so the high-level textual descriptions used by MLLM-EVAD are not sufficient on their own to detect all of the anomalies. We show that our textual descriptions do add an important new attribute, however, since, when combined with the more fine-grained attributes of Tracklet EVAL, the results with the MLLM textual descriptions improve accuracy for both datasets across all three evaluation criteria.
As shown in Table 4, on the Avenue dataset, the Tracklet-EVAL + MLLM-EVAD method improves the state-of-the-art for both the RBDC and TBDC criteria. On Street Scene (results shown in Table 3), the state-of-the-art is surpassed for the TBDC criterion. Since the frame-level criterion does not measure spatial localization accuracy which is the focus of these datasets, we would argue that the slightly-below SOTA frame-level numbers are not representative of the true accuracy of these methods. We should note that we were not able to exactly reproduce the results reported for Tracklet EVAL (Singh et al., 2024) on the Avenue dataset, most likely due to different parameters, so we report the results we could reproduce for that method and use that version in combination with MLLM-EVAD for the results on the combined method. All quantitative results reported in Tables 2, 4, and 3 are computed using raw, unsmoothed anomaly scores. Although temporal smoothing (e.g., Gaussian filtering) is commonly used for frame-level AUC evaluation, we do not apply smoothing in our main comparisons to ensure consistency across methods. For completeness, we additionally report Gaussian-smoothed frame-level results for our method on ComplexVAD in the appendix as a supplementary post-processing analysis.
@@ 4.2 Qualitative Results: Explainability
Method Explanation Quality ($\uparrow$)
Human-written $4.2\pm 0.7$
MLLM-EVAD $3.8\pm 1.1$
Table 5: Results of the human evaluation of explainability. Scores are reported as mean $\pm$ standard deviation of 5-point Likert ratings, aggregated across evaluators and anomaly clips.
In addition to the quantitative results, we present qualitative results to demonstrate the explainability capabilities of our method. The combination of using an MLLM agent and an exemplar selection strategy enables intuitive explanations for detected anomalies by providing descriptions of both the anomaly and the most similar event from the nominal training video. In Figure 2, we demonstrate four detected anomalies from ComplexVAD along with their textual descriptions and the textual descriptions of the closest exemplars, as generated by our method. For example, one detected anomaly is described as “The person is crouching down on the ground.” The closest matching exemplar from the nominal set is “The person is walking across the grass.” The action of crouching deviates from the typical behavior observed in the training data. This difference highlights how our method leverages semantic understanding to identify uncommon activity, offering an interpretable explanation for the anomaly.
Another example involves an anomaly described as “The person is holding a phone up and appearing to take a picture or video while being pushed in a large box by another person.” The closest matching exemplar from the nominal set is “The person is walking along the sidewalk.”, which is, in fact, the most common activity involving a person in this dataset. The anomalous event, a person being pushed in a box, does not occur in the nominal video and represents an unusual and complex behavior. This example further demonstrates our method’s ability to detect rare interaction patterns that deviate from typical activities in the scene. In addition, these examples illustrate how our method also provides reasonable explanations for detected anomalies. By leveraging descriptive comparisons with nominal exemplars, our method enables enhanced and explainable video anomaly detection.
As discussed in Section 5, the absence of ground-truth textual benchmarks makes quantitative evaluation of explainability difficult in the semi-supervised, single-scene VAD setting. To complement our qualitative analysis under this constraint, we perform a small-scale human evaluation of the generated explanations. We randomly select 10 anomalous video clips from ComplexVAD and, for each clip, consider two textual explanations: (i) a human-written annotation and (ii) the description generated by our system. For each clip, evaluators are shown the video and the two anonymized descriptions (labeled as “Description A” and “Description B”) presented in randomized order, without revealing their source. Evaluators rate each description on a 5-point Likert scale according to the following question: “How well does this description explain why the event in the video is anomalous?” (1: not at all informative, 5: extremely informative). We collect ratings from $5$ independent evaluators and report the mean and standard deviation of the scores aggregated across clips and evaluators in table 5. The results show that the explanations generated by MLLM-EVAD are consistently rated as informative and closely aligned with human-written explanations, supporting the interpretability claims of our approach. All human-written and MLLM-generated explanations used in this study are provided in Appendix D for qualitative inspection.
@@ 4.3 Ablation Study
We conduct two ablation studies by modifying the default settings of our method. In the first ablation study, instead of using sentence embeddings and cosine dissimilarity as the distance metric, we use the raw sentences and compute distances using BLEU (Papineni et al., 2002) and METEOR (Banerjee and Lavie, 2005). We emphasize that BLEU and METEOR are used here solely as alternative sentence similarity functions within the anomaly detection pipeline; no ground-truth textual annotations are involved, and evaluation is performed using standard VAD metrics. In Table 6, we compare the results of cosine distance over Sentence-BERT embeddings with those obtained using BLEU and METEOR, based on the descriptions generated by our method using GPT-4o as the MLLM agent. The results show that Sentence-BERT embeddings yield slightly higher accuracy in the RBDC and frame-level criteria while BLEU and METEOR perform slightly better on the TBDC criterion. Sentence-BERT is the preferred similarity function considering its greater efficiency due to embeddings which can be saved and only require a dot product for comparison.
@@ 5 Limitations and Future Directions
While our framework shows promising results, we identify several limitations that pave the way for future research. First, our implementation relies on powerful but computationally expensive multimodal models like Gemma 3. Their high inference latency poses a challenge for real-time applications. Recent studies have shown that smaller, task-specific models can perform exceptionally well on specialized tasks (Wang et al., 2020; Schick and Schütze, 2020; Marafioti et al., 2025; Mumcu and Yilmaz, 2024). We will explore fine-tuning such models for single scenes, specifically for describing object activities and interactions.
Second, a primary contribution of our method is its semantic explainability, but quantitatively evaluating this feature is difficult due to a lack of suitable benchmarks. Current datasets with textual ground truth are designed for multi-scene or weakly-supervised tasks and do not fit the semi-supervised, single-scene VAD problem (Zhu et al., 2024; Du et al., 2024; Zhang et al., 2025). Therefore, a key future direction is the development of new datasets with ground-truth textual descriptions for scene-specific anomalies.
Our framework relies on relative consistency across nominal textual descriptions rather than absolute textual correctness. Anomaly detection is performed by comparing sentence embeddings to a set of nominal exemplars, such that deviations from the learned nominal distribution determine anomaly scores. As a result, minor variations in wording or phrasing do not directly affect detection decisions, provided that nominal descriptions remain semantically consistent. In addition, we employ a fixed system prompt and a constrained task formulation that focuses on describing object activities and interactions, which further reduces sensitivity to prompt wording. Nevertheless, a more systematic evaluation of prompt sensitivity remains an important direction for future work, particularly given the computational cost of large-scale MLLM querying.
Finally, our framework currently leverages a robust traditional object detector, which is well-suited for scenarios with known object classes. To further generalize our approach, a promising future direction is the integration of open-vocabulary object detection methods (Liu et al., 2024; Cheng et al., 2024; Mumcu et al., 2025a). By enabling the recognition of an open set of object categories, this advancement would significantly broaden our system’s applicability, allowing it to analyze and describe a more diverse range of anomalies in unconstrained, real-world environments.
@@ 6 Conclusions
We demonstrate that our method outperforms existing approaches on ComplexVAD which is a challenging benchmark specifically designed to test interaction-based anomalies. In addition our method also improves of the state-of-the-art results on standard datasets like Avenue and Street Scene when combined with the fine-grained method Tracklet EVAL (Singh et al., 2024). Through ablation studies, we showed the importance of using sentence embeddings and detailed MLLM-generated descriptions. Notably, we found that Gemma 3 provided more informative descriptions than GPT-4o, contributing to higher detection accuracy.
@@
7 Broader Impact Statement
This work studies video anomaly detection with a focus on modeling scene-specific object interactions using multimodal large language models. While the proposed method is intended as a research contribution to representation learning and explainable anomaly detection, it is important to consider potential broader impacts and risks associated with its use.
A primary application domain of video anomaly detection is surveillance, which raises ethical concerns related to privacy, misuse, and surveillance overreach. If deployed irresponsibly, anomaly detection systems may contribute to unwarranted monitoring or automated enforcement. We emphasize that the proposed approach is intended as a decision-support tool rather than an automated decision-making system, and should be used in human-in-the-loop settings where outputs are interpreted by responsible operators rather than acted upon autonomously.
The use of multimodal large language models introduces additional considerations. MLLMs are computationally expensive, and converting large volumes of video data into textual descriptions increases energy consumption and carbon footprint compared to conventional VAD pipelines. For this reason, our method is not designed for real-time or large-scale deployment, and the MLLM component is used sparingly (offline and at sparse temporal intervals). Future work should consider more efficient models or distillation strategies to reduce environmental impact.
MLLM-generated textual descriptions may also hallucinate or include spurious details. In safety-critical scenarios, such hallucinations could lead to misleading explanations or mischaracterization of events. Our framework mitigates this risk by relying on relative consistency across nominal data and by using language-based representations as an auxiliary explanatory signal rather than as a sole decision criterion. Nevertheless, we caution against interpreting generated text as ground truth or using it without human oversight.
Finally, like many video analysis methods, our approach inherits biases present in training data and object detection models. Dataset bias may affect which behaviors are considered “normal” or “anomalous,” potentially leading to unfair or incorrect outcomes if deployed without careful consideration. We encourage future work to examine dataset diversity, bias mitigation, and domain-specific validation prior to any real-world deployment.
Appendix A Compute Resources
@@
For experiments using the locally deployed MLLM, we report inference latency measurements using the Gemma model. On a single NVIDIA A100 GPU, querying Gemma with paired object-centric crops and the associated prompt takes on average $0.58$ seconds per query, with a standard deviation of $0.16$ seconds (measured over repeated runs). This provides a reproducible lower bound on the inference cost of the MLLM component under local deployment.
@@ Appendix B Cropping Frames
$h$ $=\max\!\left(\tfrac{|y_{2}-y_{1}|}{2},\,h_{\text{min}}\right)$ (5)
@@
Appendix C Comparison of Gemma 3 and GPT-4o Activity Descriptions
Figure 3: The activity descriptions generated by Gemma 3 and GPT-4o with given images.
In Figure 3, we share activity descriptions generated by Gemma 3 and GPT-4o for given images and the promt: “Briefly describe what the <object name>s in the enclosed regions of these images are doing. The two images were taken one second apart.”
@@ Appendix C Comparison of Gemma 3 and GPT-4o Activity Descriptions
Appendix D Human Evaluation Examples
Clip ID Human-written Explanation MLLM-EVAD Explanation
4201 A car stops abruptly after the crosswalk, causing the car behind to halt on the crosswalk and block pedestrians before both continue. The car is driving forward as a person with a backpack walks across the crosswalk in front of it.
4254 Two people cross the crosswalk when a third runs up and jumps onto one of them, who appears distressed. The person in the black jacket appears to be pushing the person in the grey shirt, causing them to stumble forward.
4288 The skateboarder falls in the crosswalk as a scooter rider passes by. After a moment, the skateboarder gets up and leaves the crosswalk. The person is running and then falls to the ground.
4318 A dog runs across the crosswalk without a leash, moving away from its owner. The person was walking with the dog, but the dog has broken away and is now running ahead.
4360 A person hits another person on the head with a bat and the other person is visibly hurt. The two people are walking across the crosswalk together.
4371 A scooter and a bicycle collide while crossing the crosswalk, then both riders continue through. The person on the skateboard is being pulled forward by the person riding the bicycle.
4405 Two people play volleyball near the crosswalk, and their ball eventually falls toward the road. The people in the enclosed region are playing with a ball, passing it back and forth to each other.
4414 A golf cart is parked near a crosswalk and then drives away. The golf cart is driving forward as the person walks alongside it, and a car is passing in front of the golf cart.
4426 A scooter rider crosses a crosswalk, then stops, turns around, and rides back. The person is riding the scooter forward across the pavement.
4431 A person climbs a pole near the crosswalk, sits briefly, then gets down and walks away. The person is climbing onto a pole.
Table 8: Examples used in the human explainability evaluation. For each anomalous clip, we show the human-written explanation and the explanation generated by MLLM-EVAD. Anomalous clips are randomly selected from ComplexVAD dataset.
Appendix E Effect of Temporal Gaussian Smoothing
Method Frame AUC (Raw) Frame AUC (Gaussian Smoothed)
MLLM-EVAD 0.61 0.68
Scene-Graph + MLLM-EVAD 0.63 0.71
Table 9: Effect of temporal Gaussian smoothing on frame-level anomaly detection performance on the ComplexVAD dataset. Smoothing is applied only as a post-processing step for this supplementary analysis and is not used in any main quantitative comparisons.
As stated in the main text, all quantitative results reported in Tables 2, 4, and 3 are computed using raw, unsmoothed anomaly scores to ensure consistency across methods.
As a supplementary analysis, we report the effect of applying Gaussian temporal smoothing to the frame-level anomaly scores of our best-performing configuration on the ComplexVAD dataset. Temporal smoothing is a commonly used post-processing step in video anomaly detection to reduce temporal noise in frame-level scores. We emphasize that smoothing is not part of the proposed method and is not used in any of the main quantitative comparisons; the results shown here are provided solely for completeness and illustrative purposes.